Systems Staff Meeting Notes 12/27/99

Systems Staff Meeting Notes
December 27, 1999

Present: Will Chaney, Bill Covey, Debra Harris, Bobbie Meng, Everett Phillips, Suzy Shaw, Ying Tang
Announcements
  1. Hilton is stuck in traffic on the turnpike.
  2. Asbestos abatement going on over the front door of Library East. It is reportedly 'non-fryable' asbestos, meaning it is encapsulated - unable to get out into the air.
  3. Check on any old trouble tickets you may still have open & try to get them closed by the end of the year.
Review Minutes of 12/20/99
  1. RIR #2 - Spamming seems to be increasing - it has even struck the Campus Computer Coordinator List.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Received Pro-Cite upgrade for Peter McKay.
  2. Cattle still enroute.
  3. Inspiron 7500 Dell laptop received, destined to Lawan Orser.
Security
  1. Big increase in hacker network scans (including the Libraries).
  2. 'Tribal Flood' - a scan that finds ways to make someone else's computer a network bomb.
  3. Yet another Super DAT virus has been reported this week. Check the McAffee website. Bill recommends you check the site everyday.
  4. Y2K/EOY procedures - will treat this like a hurricane drill and unplug all devices, with the exception of the network & servers. Expect every failure in January to be attributed to a virus, hacker or Y2K.
Networking
  1. Bobbie put the patches on the SNA machines that were timing out. Unfortunately, most of the staff at those machines are out for the holidays and unable to give feedback.
  2. NT5 - still having problems with it not being seen from some servers.
  3. MSL feed/switch swap - need to do it at the crack of dawn or give MSL one day notice before the swap. Need to notify Pat Haskins in John Ashcraft's absence. Will also check on MSL break time hours.
  4. Network backups - having some failures on NT8. Bill suggests choosing individual items. D: drive is critical. Can change it to do just D: and hold the rest. Bill also suggests installing Backup Exec Agent on NT8, to see if it makes a difference.
  5. Not much wiring happening this week. Requests for Preservation& Norman wiring have been submitted.
  6. Novell server upgrade - 128M SIMM still not received. It can be added later, meanwhile Bill will have to download the software, then upload it to our server.
Software Projects Status
  1. Wiring database has been installed on Bobbie's machine. She will need to test it and changes will be made as necessary.
  2. IR database - Ying has finished the code to print device labels - one label is used for monitors, one for other devices. There are still some problems with the multi-user feature, need to rectify the code.
  3. Key Program database - need to enable it to save the report to text file so Facilities can send it via email.
  4. Forms to mail - Suzy needs the Linux server this week for Y2K compliance.
Calendar
  1. Rich & Winston out through today.
  2. Will out December 29-30.
  3. New Year Holiday - Friday, Dec. 31.
  4. Everett out Jan. 6-12.
  5. Debra out Jan. 20-27.
  6. Everett out Jan. 28-31.


Back to the Systems Homepage

Last updated January 3, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu